OpenStudy (anonymous):

A pipe can pump 1000 gallons of water in 4 hours. How long would it take the pipe to fill a 2500 gallon pool? A. 4 h B. 8 h C. 10 h D. 25 h

OpenStudy (mathmale):

One way to do this problem involves using an equation of ratios. 1000 gallons of water in four hours, and 2500 gallons in x hours. One way to write this equation would be\[\frac{ 1000 }{ 2500 }=\frac{ 4 }{ x }\]

OpenStudy (mathmale):

Cross multiply. then solve for x (hours).

OpenStudy (anonymous):

Do you know how to do this?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Cross multiplying gives us: 1000x=10,000

OpenStudy (mathmale):

OrangeMaster: I'd be interested in seeing how you'd explain "cross multiplication."

OpenStudy (mathmale):

WHAT YOU'RE doing is, of course, fine!

OpenStudy (anonymous):

\[1000x \div 1000=x\]\[10,000\div1000=?\]

OpenStudy (mathmale):

Sweetie: OrangeMaster hs come up with the equation 1000x=10000, and is asking you to solve it for x.

OpenStudy (anonymous):

@sweety2415, can you figure it out?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Take it this way. 1x=10 What would you multiply 1 by to get 10?

OpenStudy (anonymous):

Pump is 1000 gal/4 hours = 250 gal/hr Pool is 2500 gal Time needed is (2500 gal)/(250 gal / hr) = answer in hours

OpenStudy (anonymous):

it's to hard :(

OpenStudy (anonymous):

How many 1000's are in 10,000?

OpenStudy (anonymous):

You can even use fingers in necessary.

OpenStudy (anonymous):

Do you have Google?